# Exploiting Student Parallelism for Low-latency GPU Inference of BERT-like Models in Online Services

### v1

@@ 3.2.2. Adaptive Pruning of Student Models
To improve throughput with minimal accuracy loss, Student Parallelism further employs adaptive student pruning to reduce the student number and improve the generalization of the remained students. We follow the sub-network training in Network Architecture Searching (NAS) (cai2019once, ) and DynaBERT (hou2020dynabert, ) to prune the redundant students in boosting ensemble for final representation extraction. Then it trains the output classifier on the final representations extracted from different numbers of students to mimic the prediction output of the teacher. Since the newly added students of boosting ensemble gradually improve accuracy by reducing the residual error, we drop and prune last-trained students to reduce overall computation for higher throughput with minimal accuracy loss. Moreover, pruning can even improve accuracy and generalization by only keeping the suitable number of students, because Occam’s razor recommends searching for hypotheses constructed with the smallest possible set of elements. In § 6, we verify the reduced accuracy loss of different numbers of students and better generalization of the best-pruned set of students.
@@
Specifically, we state the details about adaptive student pruning in Algorithm 2. During the training of one data batch, It dynamically samples different numbers of students to train simultaneously (line 4). In this stage, different numbers of ensembled students generate the final representation (line 5), which has quite limited communication costs thanks to the small size of the pooled final representation. Then we put a new classifier layer on the sum of the selected students’ outputs (line 6). And we take the soft labels generated by the teacher model (line 3) as the target in the soft cross-entropy loss rather than the hard labels of the ground truth (line 7). Since there are multiple forward and back propagations of different numbers of students for the same training batch, it accumulates all the gradients to update the parameters of the classifier and all students (line 8 and 10). Finally, we evaluate different numbers of previous $k$ students on the validation set to choose the optimal student number for the best pruned students, instead of keeping all students (line 12-15).
3.2.3. Training Cost Discussion
Our offline training phase has a similar time cost to other model compressions. Although we need to train multiple students, the student is a much smaller pre-trained model compared with other baselines. And each student is only trained on the sub-sampled dataset based on the residual error in the boost ensemble learning. We can also use distributed training on multiple GPUs to speed up. Generally, Student Parallelism has a similar level of offline training cost with previous distillation works (about the 1.35x training time of DynaBERT). And the training cost can be amortized during the long-term deployment.
@@ 3.3. Online Inference
Figure 4. The online inference system designs for student parallelism: model allocation of parallel student groups and inference sample dispatching.
For efficient GPU inference for stochastic online workloads. we make specialized system designs to realize student parallelism, including student allocation and sample dispatching. We assume that the GPU cluster running online inference has multiple nodes, and each node has multiple GPUs. In general, we have one global dispatcher to balance the workloads among all nodes. And every node runs its own resource management and task scheduling. As shown in Figure 4, Student Parallelism has the following procedures during the online inference phase:
  1. (1)
  2. (2)
    User submits a request triggering one or more inference samples.
  3. (3)
  4. (4)
  5. (5)
  6. (6)
  7. (7)
    It returns results to users.
@@ 3.3.1. Student Allocation
Resource management employs a specialized student allocation strategy for the group of parallel students in every node. Since all student models share parallel and homogeneous network architecture, it can distribute the different students in the same group over the multiple GPUs in the same node as shown in Figure 4, which can better parallelize the inference of every sample. Although the parallel students in the same group need communication to gather the final representation, it only takes about 0.2 ms due to the small size of the final pooled representation and fast NV-link or PCI-E.
Student Parallelism further replicates the student group into multiple copies and assign the multiple student groups on the same set of GPUs in the node, which run in parallel to process different individual samples respectively and concurrently. Specifically, if the size of the student group is $S$ and total GPU number is $G$, it always assign the $i$-th student of $j$-th student group on the specific GPU as follows:
@@
Therefore, if $i+j\times S>G$, the i-th student in the j-th student group is co-located on the same GPU with other students in the previous groups. Student Parallelism employs NVIDIA Multi-Process Service (MPS) (mps, ) to make multiple students from different groups share the same GPU without interference. It can guarantee different students on the same GPU have separate computation units and memory bandwidth to run concurrently.
@@
During the inference, resource management keeps tracking the intensity of dynamic online workloads(i.e., the request per second) to adjust the student number accordingly. It temporarily decrease the student number (§ 3.2.2) to run more student groups for workload burst, instead of reserving some idle GPUs in advance. And after workload bursts, Student Parallelism can increase the student number to fully utilize all GPUs for high accuracy. In the initialization, resource management conduct profiling to determine the proper hyper-parameters of student allocation, e.g., the optimal student number sharing the same GPU (more details in the Appendix).
@@ 3.3.2. Inference Sample Dispatching
To reduce the extra costs of idle waiting and paddings for batching from online workloads, Student Parallelism enable the immediate and concurrent inference on different individual samples. If there is any idle student group and no buffered samples, the dispatcher can immediately send out the arrived sample for inference without waiting for batching. Since different individual samples do not arrive at the same time, different student groups can concurrently process them by overlapping between data transferring to GPU and inference computation as shown in Figure 5. Only if there are massive concurrent samples occupying all student groups, the length-aware buffer temporarily stores the new samples, which merges the samples in the same length bin into small data batches with a few paddings. To efficiently merge samples on the fly, the length-aware buffer combines the queue with a dictionary whose key-value pair is length bin and queue position respectively (details in the Appendix).
@@ 4. Evaluation
  • Student Parallelism can effectively reduce more model depth, while maintaining the best accuracy among small models without full depth or width. And adaptive student pruning can achieve comparable accuracy with even only one student pruned from all students.
@@ 4.1. Experiment Settings
4.1.1. Implementation Detials
In training, we consider all 24 compact BERT variants (turc2019well, ) and all DEBERTA variants as the student candidates to find the shallowest one for different teachers to reach target accuracy. In practice, we use the BERT-2L256D (turc2019well, ) for BERT-base, BERT-6L768D for BERT-large, and Deberta-12L768D for DEBERTA-xlarge as the backbone network and initialization of student models respectively. Following previous works (jiao2019tinybert, ; DBLP:conf/acl/XinTLYL20, ), we conduct the same data augmentation on the down-stream datasets. We conduct a grid search on various training hyperparameters to find the best for every dataset, including epoch number, batch size, learning rate, $\lambda$ weights, and data sampling ratio. And in practice, we optimize the $\alpha_{i}$ together with the student, since they are both differentiable. We run inference on 4 servers each with 4 Geforce 3090 GPUs and more implementation details about online inference systems can be found in Appendix.
@@ 4.1.2. Datasets and Metrics
We follow previous works (DBLP:conf/acl/XinTLYL20, ; sanh2019distilbert, ; hou2020dynabert, ) to use the famous GLUE benchmark (wang2018glue, ). It consists of 8 text datasets in various text mining (SST-2, CoLA, MNLI, and RTE) and information retrieval (MRPC, QQP, STS-B, and QNLI) tasks. We report various prediction quality metrics recommended by GLUE on the validation set of different tasks, including accuracy, F1 score, Matthews Correlation Coefficient (MCC) and Pearson correlation coefficient (PCC). Additionally, we also build our workload generator upon the real production trace of twitter (khaleq2018cloud, ) to feed the inference system. For discriminative text mining, we report the average latency and the 95% tail latency. And we report throughput for information retrieval tasks to better show how long users will wait for the ranked results.
@@ 4.1.3. Baselines
We compare our Student Parallelism with various types of baselines related to BERT-like models. First, we consider the finetuned 12-layer BERT-base-12L768D (devlin2018bert, ), 24-layer BERT-large-24L1024D (devlin2018bert, ), and 48-layer DeBERTa-xlarge-48L1024D (he2021deberta, ) as the original teachers. Then we conduct quantization to get I-BERT-12L768D with the same model architecture. We train the 2-layer, 6-layer, and 12-layer versions of Student Parallelism for these teachers respectively, reducing over 75% model depth. And we make the DeeBERT-12L768D (DBLP:conf/acl/XinTLYL20, ) the baseline for the adaptive computing. We use the 3-layer PKD-BERT-3L768D (sun2019patient, ), 4-layer TinyBERT-4L312D (jiao2019tinybert, ), and 6-layer DistillBERT-6L768D (sanh2019distilbert, ) as the representatives of knowledge distillation. For pruning, we choose the fastest version of DynaBERT-6L192D (hou2020dynabert, ) satisfying the target prediction. We also bring in extremely shallow finetuned models like BERT-2L256D (turc2019well, ) as the vanilla 2-layer model. However, all these baseline works can only optimize finetuned 12-layer BERT-base model. We further compare our work with Cocktail based on bagging ensemble of all the above-mentioned models except our Student Parallelism to approach the performance of 48-layer DeBERTa-xlarge. By following the previous works (zhang2019mark, ; gunasekaran2022cocktail, ), all baselines employ the dynamical batch size and ragged batch to realize adaptive batching for practical online workloads. We set the overall time budget as 10ms to dynamically set the batch size and waiting window according to profilings and online workloads (zhang2019mark, ).
@@ 4.2. Overall Performance
In this subsection, we first verify if Student Parallelism can maintain the accuracy. Then we show its advantages in inference efficiency, including average latency, tail latency and throughput.
@@ 4.2.1. Prediction Quality
As shown in Table 1, Student Parallelism can achieve the best overall score up to 96.32% of BERT-base with only two layers, and it is the only one compressing BERT-large and DeBERTa-xlarge into their one-quarter model depth while reaching the target accuracy. The results show other baselines can only reduce about half the layers of the BERT-base to maintain the accuracy, which reach the target score (95% overall score of the teachers, as recommended in MLPerf inference benchmark (reddi2020mlperf, )). Compared with tiny baselines with less than 6 layers, Student Parallelism has significant advantages in the prediction qualities. Especially, the 3-layer PKD-BERT and 2-layer BERT-2L512D can only maintain less than 90% overall score. Student Parallelism is also better than any 6-layer or deeper reduced baselines with over 3$\times$ larger model depth. It is even comparable with I-BERT always having full depth and width. Because of improved generalization from adaptive student pruning, our work can use even one student to handle workload burst and reach a competitive overall score (95.25% of the BERT-base) compared with 6-layer baselines. Additionally, if using deeper student models, Student Parallelism also can reduce the model depth of larger BERT-large and DeBERTa-xlarge by 75%, while the others fail. And Cocktail cannot reach the target of DeBERTa-xlarge even with 24-layer BERT-large. Since DeeBert and Cocktail rely on the classification probability to work, we omit their results in the regression dataset STS-B.
@@ 4.2.3. Throughput
In information retrieval tasks, like the search engine and QA system, the user has to wait to get the final result until dozens of document candidates are scored for the relevance of the query. Hence, we report the throughput per GPU for best-pruned students and the single student to compare with other baselines on the same number of GPUs. Because some models have been out of memory, we set the batch size as 64 for all baselines. The last column of Table 2 records the overall throughput / GPU number for all methods in the MRPC dataset. Our work can improve throughput over the others by up to 22.27×22.27$\times$22.27 ×, because it can leverage adaptive student pruning to use only one small student whose generalization has been improved (§ 4.2.1). Meanwhile, even the best pruned of students (3 in total) can achieve comparable throughput with some competitive baselines like TinyBERT for better accuracy if the workload is not so heavy. Except for the embedding layer, all model parameters mean computations, so the throughput negatively correlates with the model size. Even with fast integer calculation, the full model size of I-BERT limits its throughput improvement to 3×3$\times$3 ×. The 6-layer DynaBERT is the most competitive one of the small baselines reaching the target accuracy, because it reduces model width to be $25\%$. The vanilla BERT-2L256D has high throughout, but it is far away from the target accuracy. And Cocktail suffers from low throughput per GPU since the bagging ensemble has to use multiple GPUs.
@@ 4.3. Abalation Studies
In this section, we further conduct ablation studies in Student Parallelism, including hyperparameters tuning, accuracy analysis experiment, and latency analysis experiment.
@@ 4.3.1. Model Hyperparameter Tuning
(a) Student Number
(b) Model Depth
Since student number and model depth are key hyperparameters influencing the prediction quality, we conduct experiments to measure how they affect the prediction quality. According to the results, we verify that adaptive student pruning can mprove the generalization of the remained students and reduce the student number with minimal accuracy loss. Furthermore, we can even further improve prediction quality with larger model depth with the relaxed latency requirement.
@@
Figure 6(a) shows that adaptive student pruning consistently outperforms the direct boosting ensemble of the same number of students without pruning in overall GLUE scores. The results verify that adaptive student pruning can effectively improve the generalization as a pruning technology. Because it improves the generalization of the remained student, 3 students are the best pruned to get the highest test performance in the most datasets. And even only one student pruned from all 6 students can maintain a relatively high score. However, when we directly train the boosting ensemble of $m$ students without pruning, the limited total student number has more difficulties in finding the optimal solution.
As shown in Figure 6(b), our Student Parallelism can further improve the prediction with a deeper student when learning from BERT-base in all datasets. The 4-layer students distilled from BERT-base can significantly improve the prediction results. Student Parallelism can even get a similar F1 score to the original 12-layer BERT-base if using 6-layer students. And compared with other best-optimized baselines that have the same layers (such as BERT-2L256D, TinyBERT, and DistillBERT), our Student Parallelism can consistently outperform them in overall GLUE scores, while having the same layer number and similar inference latency.

### v2

@@ 3.2. Adaptive Student Pruning
By conducting adaptive student pruning, we further train the classifier layer on top of the above final representation for better generalization and dynamical student numbers. We follow the sub-network training in Network Architecture Searching (NAS) (cai2019once, ) to dynamically prune the redundant students for final representation extraction. Then it trains the output classifier on the final representations extracted from different numbers of students to mimic the prediction output of the teacher.Since the newly added students of boosting ensemble gradually improve accuracy by reducing the residual error, we drop and prune last-trained students without computing the importance scores in previous pruning works (hou2020dynabert, ). Because of Occam’s razor theory, student pruning can improve accuracy and generalization by removing the redundant students and only keeping the suitable group size. Moreover, adaptive student pruning enables dynamic student numbers with minimal accuracy loss. For high throughput, we drop any numbers of the tailing students, while the remained students are fine-tuned in this setting to get higher accuracy. § 4.3.1 shows the better generalization of the best-pruned set of students and reduced accuracy loss of different numbers of students.
@@
Specifically, we state the details about adaptive student pruning in Algorithm 2. During the training of one data batch, It dynamically samples different numbers of students to train simultaneously (line 4 in Algorithm 2). In this stage, different numbers of ensembled students generate the final representation (line 5), which has quite limited communication costs thanks to the small size of the pooled final representation. Then we put a new classifier layer on the sum of the selected students’ outputs (line 6). We take the soft labels generated by the teacher model (line 3) as the target in the soft cross-entropy loss rather than the hard labels of the ground truth (line 7). Since there are multiple forward and back propagations of different numbers of students for the same training batch, it accumulates all the gradients to update the parameters of the classifier and all students (line 8 and 10). Finally, we evaluate different numbers of previous $k$ students on the validation set to choose the optimal student number for the best-pruned students, instead of keeping all students (line 12-15).
@@ 3.3. Online Inference
Figure 4. The online inference overview and procedure of student parallelism: student allocation and task dispatching
For efficient GPU inference on online workloads, we make specialized student allocation and inference sample dispatching for student parallelism. Therefore, we can reduce the extra costs of waiting and padding to collect large and regular data batches for data parallelism. We assume that the GPU cluster running online inference has multiple nodes, and each node has multiple GPUs. And we have one global dispatcher to balance the workloads among all nodes. Figure 4 shows the overview and procedure of online inference(more details in Appendix B.4).
@@ 3.3.1. Student Allocation
As shown in Figure 4, Student Parallelism can distribute the different students over the multiple GPUs in the same node, which can better parallelize the inference of every sample. Although the parallel students in the same group need communication to gather the final representation, it only takes about 0.2 ms due to the small size of the final pooled representation and fast NV-link or PCI-E. It further replicates the student group into multiple copies that are running on the same set of GPUs in the node. Specifically, if the size of the student group is $S$ and total GPU number is $G$, the $i$-th student of $j$-th student group is allocated as follows:
@@
Therefore, if $i+j\times S>G$, the student co-locates on the same GPU with other students in the previous groups, which run in parallel to process different individual samples concurrently.
During the inference, it keeps tracking the intensity of dynamic online workloads(i.e., the request per second) to adjust the student number accordingly. It temporarily decreases the student number (§ 3.2) to run more student groups for workload bursts, instead of reserving some idle GPUs in advance. After workload bursts, Student Parallelism can increase the student number to fully utilize all GPUs for high accuracy. More details about MPS and allocation hyper-parameters can be found in Appendix B.4.1).
@@ 3.3.2. Inference Sample Dispatching
Figure 5. Length-aware Buffer: New sequence sample (e.g., the red one whose length is 30) can be merged with the previous element into a small batch, since they share the same bin and the element size is not full. Otherwise, the new sequence sample (e.g., the blue one of 52 tokens cannot join in the full element) is added to the buffer tail as a new element.
To reduce the extra costs of idle waiting and paddings for batching from online workloads, Student Parallelism conduct the immediate and concurrent inference on different individual samples. If there is any idle student group and no buffered samples, the dispatcher can immediately send out the arrived sample for inference without waiting for batching. Since different individual samples do not arrive at the same time, different student groups can concurrently process them by overlapping between data transferring to GPU and inference computation. Only if there are massive concurrent samples occupying all student groups, the length-aware buffer temporarily stores the new samples, which merges the samples in the same length bin into small data batches with a few paddings. To efficiently merge samples on the fly, the length-aware buffer combines the queue with a dictionary whose key-value pair is length bin and queue position respectively to achieve $O(1)$ time complexity, as shown in Figure 5. More details about sample dispatching based on length-aware buffer can be found in Appendix B.4.2.
3.3.3. Efficiency Analysis
We build a performance model to make a quantitative efficiency analysis. By considering all model and data factors like model depth $D$, model width $W$, batch size $B$, and input length $N$, waiting time $Q(B)$, parallel computation capability $C$, the PCI-E transferring bandwidth $T$ , the GPU number $G$, and the total parallel model number $M$, we can model the latency as follows:
(7) $\text{Latency}=\Bigl{\lceil}\frac{WBN^{2}M}{CG}\Bigr{\rceil}+Q(B)+\frac{BN}{T}\$
Obviously, Student Parallelism has the advantage of low model depth and is free of waiting and padding, compared to the existing works. Please refer to Appendix A.2 for more details about the quantitative analysis and comparison.
@@ 4. Evaluation
  • Student Parallelism can effectively reduce more model depth while maintaining the best accuracy among small models without full depth or width. Adaptive student pruning can achieve comparable accuracy with even only one student pruned from all students.
@@ 4.1. Experiment Settings
4.1.1. Implementation Details
In training, we consider all 24 compact BERT variants (turc2019well, ) and all DEBERTA variants as the student candidates to find the shallowest one for different teachers to reach target accuracy. In practice, we use the BERT-2L256D (turc2019well, ) for BERT-base, BERT-6L768D for BERT-large, and Deberta-12L768D for DEBERTA-xlarge as the initialization of student models. We conduct the same data augmentation with previous works (jiao2019tinybert, ; DBLP:conf/acl/XinTLYL20, ). We conduct a grid search on various training hyperparameters to find the best for every dataset, including epoch number, batch size, learning rate, $\lambda$ weights, and data sampling ratio. And in practice, we optimize the $\alpha_{i}$ together with the student, since they are both differentiable. We use 16 Geforce 3090 GPUs on 4 nodes and more implementation details can be found in Appendix B.
@@ 4.1.2. Datasets and Metrics
We follow previous works (DBLP:conf/acl/XinTLYL20, ; sanh2019distilbert, ; hou2020dynabert, ) to use the famous GLUE benchmark (wang2018glue, ). It consists of 8 text datasets in various text mining (SST-2, CoLA, MNLI, and RTE) and information retrieval (MRPC, QQP, STS-B, and QNLI) tasks. We report various prediction quality metrics recommended by GLUE, including accuracy, F1 score, Matthews Correlation Coefficient (MCC), and Pearson correlation coefficient (PCC). Additionally, we also build our workload generator upon the real production trace of twitter (khaleq2018cloud, ) to feed the inference system. For text mining, we report the average latency and the 95% tail latency. We report throughput for information retrieval tasks to better show how long users will wait for the ranked results.
@@ 4.1.3. Baselines
We compare our Student Parallelism with various types of baselines related to BERT-like models. First, we consider the finetuned 12-layer BERT-base-12L768D (devlin2018bert, ), 24-layer BERT-large-24L1024D (devlin2018bert, ), and 48-layer DeBERTa-xlarge-48L1024D (he2021deberta, ) as the original teachers. Then we conduct quantization to get I-BERT-12L768D with the same model architecture. We train the 2-layer, 6-layer, and 12-layer versions of Student Parallelism for these teachers respectively, reducing over 75% model depth. And we make the DeeBERT-12L768D (DBLP:conf/acl/XinTLYL20, ) the baseline for the adaptive computing. We use the 3-layer PKD-BERT-3L768D (sun2019patient, ), 4-layer TinyBERT-4L312D (jiao2019tinybert, ), and 6-layer Distilbert-6L768D (sanh2019distilbert, ) as the representatives of knowledge distillation. For pruning, we choose the fastest version of DynaBERT-6L192D (hou2020dynabert, ) satisfying the target prediction. We also bring in extremely shallow finetuned models like BERT-2L256D (turc2019well, ) as the vanilla 2-layer model. However, all these baseline works can only optimize the finetuned 12-layer BERT-base model. We further compare our work with Cocktail based on bagging ensemble of all the above-mentioned models except our Student Parallelism to approach the performance of 48-layer DeBERTa-xlarge. By following the previous works (zhang2019mark, ; gunasekaran2022cocktail, ), all baselines employ the dynamical batch size for practical online workloads. We set the time budget as 10ms to dynamically set the batch size and waiting window according to online workloads (zhang2019mark, ).
@@ 4.2. Overall Performance
In this subsection, we first verify if Student Parallelism can maintain the accuracy. Then we show its advantages in average latency, tail latency and throughput.
@@ 4.2.1. Prediction Quality
As shown in Table 1, Student Parallelism can achieve the best overall score up to 96.32% of BERT-base with only two layers, and it is the only one compressing BERT-large and DeBERTa-xlarge into their 25% model depth while reaching the target accuracy. The results show other baselines can only reduce about half the layers of the BERT-base to maintain accuracy, which reach the target score11195% overall score of the teachers, recommended by MLPerf (reddi2020mlperf, ). Compared with tiny baselines with less than 6 layers, Student Parallelism has significant advantages in the prediction qualities. Especially, the 3-layer PKD-BERT and 2-layer BERT-2L512D can only maintain less than 90% overall score. Student Parallelism is also better than any 6-layer or deeper reduced baselines with over 3$\times$ larger model depth. It is even comparable with I-BERT always having full depth and width. Because of improved generalization from adaptive student pruning, our work can use even one student to handle workload burst and reach a competitive overall score (95.25% of the BERT-base) compared with 6-layer baselines. If using deeper student models, Student Parallelism also can reduce the model depth of larger BERT-large and DeBERTa-xlarge by 75%, while the others fail. Cocktail cannot catch up with DeBERTa-xlarge even with 24 layers. Because DeeBert and Cocktail need classification probability to work, they cannot handle regression dataset STS-B.
@@ 4.2.3. Throughput
In information retrieval tasks, like the search engine and QA system, the user has to wait to get the final result until dozens of document candidates are scored for the relevance of the query. Hence, we report the throughput per GPU for best-pruned students and the single student to compare with other baselines on the same number of GPUs. Because some models have been out of memory, we set the batch size as 64 for all baselines. The last column of Table 2 records the overall throughput / GPU number for all methods in the MRPC dataset. Our work can improve throughput over the others by up to 22.27×22.27$\times$22.27 ×, because it can leverage adaptive student pruning to use only one small student whose generalization has been improved (§ 4.2.1). Meanwhile, even the best pruned of students (3 in total) can achieve comparable throughput with some competitive baselines like TinyBERT for better accuracy if the workload is not so heavy. Except for the embedding layer, all model parameters mean computations, so the throughput negatively correlates with the model size. Even with fast integer calculation, the full model size of I-BERT limits its throughput improvement to 3×3$\times$3 ×. The 6-layer DynaBERT is the most competitive one of the small baselines reaching the target accuracy because it reduces model width to be $25\%$. The vanilla BERT-2L256D has high throughout, but it is far away from the target accuracy. Cocktail suffers from low throughput per GPU since it has to use multiple GPUs.
@@ 4.3. Abalation Studies
We further conduct ablation studies, including hyperparameters tuning, accuracy analysis, and latency analysis.
@@ 4.3.1. Model Hyperparameter Tuning
Since student number and model depth are key hyperparameters influencing the prediction quality, we conduct experiments to measure how they affect the prediction quality. According to the results, we verify that adaptive student pruning can improve the generalization of the remained students and reduce the student number with minimal accuracy loss. Furthermore, we can even further improve prediction quality with larger model depth with the relaxed latency requirement.
@@
Figure 6(a) shows that adaptive student pruning consistently outperforms the direct boosting ensemble of the same number of students without pruning in overall GLUE scores. The results verify that adaptive student pruning can effectively improve the generalization as a pruning technology. Because it improves the generalization of the remained students, 3 students are the best pruned to get the highest test performance in most datasets. And even only one student pruned from all 6 students can maintain a relatively high score. However, when we directly train the boosting ensemble of $m$ students without pruning, the limited total student number has more difficulties in finding the optimal solution.
As shown in Figure 6(b), our Student Parallelism can further improve the prediction with a deeper student when learning from BERT-base in all datasets. The 4-layer students distilled from BERT-base can significantly improve the prediction results. Student Parallelism can even get a similar F1 score to the original 12-layer BERT-base if using 6-layer students. And compared with other best-optimized baselines that have the same layers (such as BERT-2L256D, TinyBERT, and Distilbert), our Student Parallelism can consistently outperform them in overall GLUE scores, while having the same layer number and similar inference latency.
(a) Student Number
(b) Model Depth